Two circles have the same center. The radius of the larger circle is 3 units longer than the radius of the smaller circle. Find the difference in the circumfrences of the two circles. Round to the nerest hundredth.
Answer by Alan3354(21583)

Two circles have the same center. The radius of the larger circle is 3 units longer than the radius of the smaller circle. Find the difference in the circumfrences of the two circles. Round to the nerest hundredth.
------------------
C = 2*pi*r
C2 = 2*pi*(r+3) = 2*pi*r + 6*pi
diff = 6pi = ~18.85 units
--------------
Having the same center is not relevant.